How many grams of HCl are produced according to the following equation,

H2 + Cl2 → 2 HCl, when 4.0 g of hydrogen reacts completely?

with work and explanations, please using conversion factors
Answer #1


Molar mass of H2 = 2.016 g/mol

mass of H2 = 4 g
mol of H2 = (mass)/(molar mass)
= 4/2.016
= 1.984 mol


Balanced chemical equation is:
H2 + Cl2 → 2 HCl


According to balanced equation
mol of HCl formed = (2/1)* moles of H2
= (2/1)*1.984
= 3.968 mol


Molar mass of HCl,
MM = 1*MM(H) + 1*MM(Cl)
= 1*1.008 + 1*35.45
= 36.458 g/mol


mass of HCl = number of mol * molar mass
= 3.968*36.46
= 145 g
Answer: 145 g